OpenStudy (anonymous):

Help !! Giving Out Medals ;) Evaluate 5a + 5b , where a= -6 and b= -5

OpenStudy (texaschic101):

This is just a matter of subbing then solving 5a + 5b when a = -6 and b = -5 5(-6) + 5(-5) = -30 + (-25) = -30 - 25 = - 55